# Kinetic formulation of compartmental epidemic models

### v1

@@ Abstract
We introduce a kinetic model that couples the movement of a population of individuals with the dynamics of a pathogen in the same population. The model is formally shown to generalize the well-known compartmental models in mathematical epidemiology. We prove the existence and uniqueness of solutions in appropriate spaces for particular instances of the model. We finish with some examples, and discuss possible applications of this modeling approach.
@@ 1 Introduction
Kinetic models were introduced in the 19th century in the study of gas dynamics [Cer69, Cer88]; in the late 20th century, they found many applications in biological modeling, from which we highlight applications in chemotaxis [Pat53, ODA88, OS97, CMPS04, CDSM+06], immunology [DMMRRS22, ORRS24], cancer dynamics [BBD04, BD04], and, more recently, epidemiology [BDP21, BBDP21].
On the other hand, mathematical models of epidemiology have a long story that dates back to seminal works by Daniel Bernoulli in the 18th Century, followed closely by Jean d’Alembert. In these works, they studied the possible impact of variolation, a primitive technique akin to vaccination (but with some important differences); cf. [Bac11] for a detailed historical account with modern mathematical notation. After a long gap, in the early 20th Century, what is now known as compartmental models gained maturity. We refer to [Ros16, KMW27] for the original references, and [Thi03, Bri03, Mur02] for a modern approach.
In the present work, we consider the so-called SIRS model, in which the population is divided into three classes: the Susceptible, individuals that can be contaminated by a certain pathogen, the Infectious, individuals that can transmit the pathogen, and the Recovered, individuals that, after being infected, profit from a temporary or permanent immunity from the disease; when the immunity is permanent, it is customary to refer simply to the “SIR model”.
@@
One of the most well-known realizations of the above model is through the use of ordinary differential equations (ODE), in which case it reads
@@
We start by introducing a model inspired by the kinetic theory of gases, in which individuals move in a certain two-dimensional space; the movement is composed of a random reorientation, perturbed by a drift in a certain preferred direction that possibly depends on the full distribution of individuals (e.g., to avoid clusters of infectious individuals). Furthermore, depending on the encounters in state space (i.e, the space including not only position but also displacement velocity), individuals change their state from S to I, to R, and back to S, according to the SIRS dynamics discussed before. See Fig 1 for an illustration.
Figure 1: Illustration of the kinetic models for three classes of particles moving in a two-dimensional space. Each circle represents an individual, and the arrow indicates his or her velocity vector. Movements are composed of random reorientations perturbed by a density-dependent drift. Blue individuals are of S type, while red and green individuals are of I and R types, respectively. In the highlight, we see an encounter in phase space (i.e., same position in space and same velocity) of an S and an I individual, possibly resulting in two I individuals, according to the prescribed transition rate. Note that when two individuals have the same position and velocity, there is more opportunity (i.e., time) for the transmission to happen than if only the same position were required.
@@
In [HC24], the SIRS evolutionary dynamics is written in the framework of evolutionary game theory, showing that the SIRS model can be reinterpreted as a mix of two different games, with two different group sizes.
Putting these two ideas together, we have a kinetic model with a transition kernel inspired by epidemic models.
There are several differences between the model introduced here and the one presented in [BDP21] that are worth a brief comment. The most important one is that we include class transitions directly inspired by the SIRS model, with the aim of having a macroscopic limit, when the problem is spatially homogeneous, the system (1). A second important difference is to consider more general movements of the individuals and proceed to a careful scale separation of the movement and the class transitions. Finally, we study the existence of solutions of the kinetic model, although in a limited manner. However, there are clear intersection points that could be explored in future works.
We finish this introduction with an outline of the present work. In Sec. 2 we introduce our basic model, i.e, the kinetic model for epidemiology; in Sec. 3 we formally show that its form macroscopic limit is, in the spatially homogeneous case, the SIRS model (1), validating the idea that the kinetic model is a generalization of the SIRS model. Furthermore, we define a mesoscopic model, a partial differential equation (PDE) of drift-diffusion-advection type, that lies in between the kinetic and the SIRS model. We proceed in Sec. 4 to show the existence of solutions in the appropriate spaces of particular instances of the kinetic and, by extension, of the PDE model derived before. We show three examples in Sec. 5 and discuss possible limitations and continuations of the present work in Sec. 6.
@@ 2 The Model
Let $f_{j}:\mathbb{R}_{+}\times\Omega\times V$, $j=S,I,R$, be the density of individuals of class $j$ at time $t\geq 0$, position $x\in\Omega\subset\mathbb{R}^{2}$ and velocity $v\in V\subset\mathbb{R}^{2}$. We assume that the space domain $\Omega$ is a bounded open set with smooth boundary and $V$ is compact and spherically symmetrical (in particular, $\int_{V}v\mathrm{d}v=0$). We define $|V|:=\int_{V}\mathrm{d}v$. Note that both $\Omega$ and $V$ are class-independent (i.e., do not depend on $j$); however, check Subsec. 5.3.
@@
In the sequel, whenever we refer to results valid to all classes of individuals, we use the subindex $j$, without further specifications. We use the notation $f_{j}:=f_{j}(t,x,v)$, and $f^{\prime}_{j}:=f_{j}(t,x,v^{\prime})$.
The transition kernel $\mathfrak{Q}_{j}$ describes three different behaviors for the individual in the population: i) a random movement, associated to a mean regression with respect to the equilibrium distribution; ii) a preferred direction in which individuals tends to move towards more favorable environments; and iii) a class-transition kernel accounting to the pathogen spread dynamics, healing and immunity waning, according to the SIRS model.
@@
  • The random reorientation kernel is given by
@@
    The reorientation rate is given by $\lambda_{j}>0$. This particular expression means that, given a point $x\in\Omega$, the number of individuals with direction $v$ tends to increase if it is below average and decrease otherwise. We note that the kernel $\mathcal{Q}_{j}^{(R)}$ is conservative, i.e., $\int_{V}\mathcal{Q}_{j}^{(R)}[\mathbf{f}]\mathrm{d}v=0$. This simple fact will be useful in the subsequent calculations.
@@
    $\mathcal{Q}_{S}^{(P)}[f_{I},f_{S}]$ $:=\mathcal{Q}_{S}^{(P)}[\mathbf{f}]$
    $=\int_{V}\bigl{(}T[f_{I}](v,v^{\prime})f_{S}(t,x,v^{\prime})-T[f_{I}](v^{\prime},v)f_{S}(t,x,v)\bigr{)}\mathrm{d}v^{\prime},$
    where the transition from direction $v^{\prime}$ to $v$ is given by
@@
    with $\eta,\xi\geq 0$. In particular, this means that susceptible individuals try to avoid the direction with increasing concentrations of I individuals (first term in the right-hand side) and to keep the direction if it points toward a decrease of the concentration of I individuals. For completeness, we also define preferred direction kernels for I and R classes to be identically null:
    $\mathcal{Q}_{I}^{(P)}[\mathbf{f}]:=\mathcal{Q}_{R}^{(P)}[\mathbf{f}]:=0\ ,\quad\forall\enspace\mathbf{f}\ .$
@@
The conservative property of the kernels $Q^{(R)}_{j}$ and $Q^{(P)}_{j}$ is not true in general, as class transitions are a fundamental feature of the proposed model, and the kernels $Q^{(T)}_{j}$ do not satisfy it. However, as we will see, the non-conservative part of the kernel is of higher order, and it will not introduce extra difficulties in the calculations of the formal limits.
@@
See Fig. 2 to an illustration with the numeric implementation of the model.
Figure 2: Representation of the mesoscopic density $\rho_{S}$ (blue, left), $\rho_{I}$ (red, middle), and $\rho_{R}$ (right, green). More intense colors indicate higher concentrations. We consider the initial condition given by S and I individuals in the center of a square lattice of size $50\times 50$, with random initial velocities. Initial conditions are far from the boundary, and for times considered, they were irrelevant. Initially, there are 10000 S individuals and 50 I individuals. We plot the configuration after 3, 6, 9, 12, and 15 steps, each step including the update of all particles. Class-transition parameters were given by $\alpha=0$ (i.e., immunity is permanent), $\beta=0.75$ (each encounter between S and I individuals has a probability of $\nicefrac{{3}}{{4}}$ to generate a new I individual), and $\gamma=0.5$ (i.e., in each time step, each I individual has a probability $\nicefrac{{1}}{{2}}$ to became R. The random reorientation parameter is $\lambda=0.5$ and we did not consider preferred directions, i.e., $Q^{(P)}_{j}\equiv 0$.
@@ 3 The macroscopic limit
In this section, we study the macroscopic limit of the kinetic model in three steps. In Subsec. 3.1 we rewrite Eq. (2) in non-dimensional form, and identify small parameters; in the sequel, Subsec. 3.2, we identify an advection-reaction-diffusion partial differential equation (PDE) for the evolution of the density of individuals in the domain $\Omega$. We conclude in Subsec. 3.3, that spatially homogeneous solutions of the previous system of PDEs, whenever they exist, correspond to the solution of the SIRS epidemic model (1).
@@ 3.1 Adimensionalization of the kinetic model
In order to provide the dimensionless form of Eq. (2), we define typical observation times, spaces, velocities, and densities $t_{0}$, $x_{0}$, $v_{0}$, $f_{0j}$, respectively. We define new adimensional variables $\tilde{t}:=t/t_{0}$, $\tilde{x}:=x/x_{0}$, $\tilde{v}:=v/v_{0}$, and $\tilde{f}_{j}:=f_{j}/f_{0j}$ such that
@@
We assume that typical density is low (an assumption akin to the rarefied gas assumption used in taking macroscopic limits of ideal gases modeled by Boltzmann equations). More precisely, we redefine the transition kernels $\varepsilon^{-2}\tilde{\mathfrak{Q}}_{j}[\mathbf{f}]=\frac{t_{0}}{f_{0j}}\mathfrak{Q}_{j}[\mathbf{f}_{0}\tilde{\mathbf{f}}]$, and omit the $\tilde{}$ notation to find the non-dimensional form of the kinetic equation
$\varepsilon^{2}\partial_{t}f_{j,\varepsilon}+\varepsilon v\cdot\nabla_{x}f_{j,\varepsilon}=\mathfrak{Q}_{j,\varepsilon}[\mathbf{f}_{\varepsilon}],$ (3)
where $\varepsilon$ is a small positive dimensionless parameter.
@@
$\mathfrak{Q}_{j,\varepsilon}=\mathcal{Q}_{j}^{(R)}+\varepsilon\mathcal{Q}_{j}^{(P)}+\varepsilon^{2}\mathcal{Q}_{j}^{(T)}\ .$
This means, in particular, that the leading order effect in the individual movement is the reorientation, followed by a small-order perturbation that gives preferred directions (when it is non-zero), and a second-order effect guaranteeing the class transitions are compatible with the SIRS model.
In the next subsection, we start to derive a partial differential equation for the purely spacial dependence of the leading order term, i.e., to
$\rho_{j}:=\lim_{\varepsilon\to 0}\int_{V}f_{j,\varepsilon}\mathrm{d}v\ .$
@@ 3.2 The mesoscopic limit
Theorem 1 (Lax-Milgram Theorem [Ram80, Sho97, BCST25]).
[…]
In the sequel, we will use a more specific form of the Lax-Milgram theorem [^3] There is a slight confusion in the notation here, as the definition of the reorientation kernel was done in boldface variables. Here, we also omit the subindex $j$. the kernel $Q^{(R)}_{j}[\mathbf{f}]$, as defined in Sec. 2, depends only on $f_{j}$, and for the proof of the corollary we write only $Q^{(R)}[f]$ for $Q^{(R)}_{j}[\mathbf{f}]$. We hope this causes no confusion.:
@@
Let $\mathcal{H}_{0}=\{f\in L^{2}(V)|\int_{V}f\mathrm{d}v=0\}$. Then, for each $g\in\mathcal{H}_{0}$ there is a unique $f\in\mathcal{H}_{0}$ such that $Q^{(R)}[f]=-g$, in the weak sense.
@@
It is clear that $\mathcal{H}_{0}$ is a Hilbert space with the natural $L^{2}$-inner product, $\langle f,g\rangle:=\int_{V}fg\mathrm{d}v$. Define the bilinear form $\mathcal{B}(g,h):=-\int_{V}Q^{(R)}[g]h\mathrm{d}v$, and
$\mathcal{B}(g,g)=-\lambda_{j}\left(\int_{V\times V}\frac{1}{|V|}g^{\prime}g\mathrm{d}v^{\prime}\mathrm{d}v-\int_{V}g^{2}\mathrm{d}v\right)=\lambda_{j}\langle g,g\rangle.$
From the fact that $\lambda_{j}$ does not depend on $v$ and is strictly positive, we conclude that $\mathcal{B}$ is coercive. The continuity is immediate.
@@
$\langle-g,h\rangle=-\mathcal{B}(f,h)=\int_{V}Q^{(R)}[f]h\mathrm{d}v\ ,\quad\forall h\in\mathcal{H}_{0}\ ,$
which finishes the proof. ∎
@@
$f_{j,\varepsilon}=f_{j,0}+\varepsilon f_{j,1}+\varepsilon^{2}f_{j,2}\ .$
Note that $f_{j,1}$, $f_{j,2}$ are not particular instances of $f_{j,\varepsilon}$.
The objective of this subsection is to find a system of PDEs for $\rho_{j}=\int_{V}f_{j,0}\mathrm{d}v$.
Remark 2.
Although the notation can be confusing at this point, we call the reader’s attention to the fact that $f_{j,0}$ is the leading order expansion of $f_{j}$ and not the density normalization $f_{0j}$ introduced before.
Expanding both sides of Eq. (3) we find
@@
$\qquad+\varepsilon^{2}\left(\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{2}]+\mathcal{Q}_{j}^{(P)}[\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(T)}[\mathbf{f}_{0}]\right)\ .$

### v2

@@ Abstract
We introduce a kinetic model that couples the movement of a population of individuals with the dynamics of a pathogen in the same population. We consider that transmission occurs when a susceptible and an infectious individual are sufficiently close for a sufficiently long time. We show that the model is formally compatible with the well-known SIRS model in mathematical epidemiology. Namely, after identifying an appropriate dimensionless variable, and considering the limit when that variable is small, we introduce a partial differential equation model of advection-drift-diffusion type (mesoscopic model), which for spatially homogeneous solutions reduces to the SIRS model. We prove the existence and uniqueness of solutions in appropriate spaces for particular instances of the model. We finish with some examples and discuss possible applications and generalisation of this modelling approach, linking kinetic models, evolutionary game theory, and mathematical epidemiology.
@@ 1 Introduction
Kinetic models were introduced in the 19th century in the study of gas dynamics [Cer69, Cer88]; in the late 20th century, many applications in biological modeling were found, from which we highlight applications in chemotaxis [Pat53, ODA88, OS97, CMPS04, CDSM+06], immunology [DMMRRS22, ORRS24], cancer dynamics [BBD04, BD04], and, more recently, epidemiology [BDP21, BBDP21].
Mathematical models of epidemiology have a long story that dates back to seminal works by Daniel Bernoulli in the 18th Century, followed closely by Jean d’Alembert. In these works, they studied the possible impact of variolation, a primitive technique akin to vaccination (but with some important differences); cf. [Bac11] for a detailed historical account with modern mathematical notation. After a long gap, in the early 20th Century, what is now known as compartmental models gained maturity. We refer to [Ros16, KMW27] for the original references, and [Thi03, Bri03, Mur02] for a modern approach.
In the present work, we consider the so-called SIRS model, in which the population is divided into three classes: the Susceptible, individuals that can be contaminated by a certain pathogen, the Infectious, individuals that can transmit the pathogen, and the Recovered, individuals that, after being infected, benefit from a temporary or permanent immunity from the disease; when the immunity is permanent, it is customary to refer simply to the “SIR model”.
@@
One of the most well-known realisations of the above model is through the use of ordinary differential equations (ODE), in which case it reads
@@
We start by introducing a model inspired by the kinetic theory of gases, in which individuals move in a certain space; the movement is composed of a regression toward the mean, perturbed by a drift in a certain preferred direction that possibly depends on the full distribution of individuals (e.g., to avoid clusters of infectious individuals). Furthermore, depending on the encounters in state space (i.e, the space including not only position but also displacement velocity), individuals change their state from S to I, to R, and back to S. However, after introducing space and velocity in the SIRS model, transitions will depend not only on the total amount of individuals in each class, as in Eqs. (1), but on their relative positions and movement. More precisely, transitions from S to I, i.e., disease transmission, will be more likely if individuals are close to each other (at the same point of space), and if their contact lasts longer (i.e., with null relative motion). See Fig 1 for an illustration of a kinetic model in two-dimensions.
Figure 1: Illustration of the kinetic models for three classes of particles moving in a two-dimensional space. Each circle represents an individual, and the arrow indicates his or her velocity vector. Movements are composed of regression toward the mean velocity perturbed by a density-dependent drift. Blue individuals are of S type, while red and green individuals are of I and R types, respectively. In the highlight, we see an encounter in phase space (i.e., same position in space and same velocity) of an S and an I individual, possibly resulting in two I individuals, according to the prescribed transition rate. Note that when two individuals have the same position and velocity, there is more opportunity (i.e., time) for the transmission to happen than if only the same position were required.
@@
In [HC24], the SIRS evolutionary dynamics was written in the framework of evolutionary game theory, showing that the SIRS model can be reinterpreted as a mix of two different games, with two different group sizes.
Putting these two ideas together, we have a kinetic model with a transition kernel inspired by epidemic models. From the kinetic model, and following classical procedures, we compute the so-called macroscopic limit. In the specialised literature, it is known as hydrodynamical limit. Namely, it consists of the derivation of fluid dynamic equations from the detailed molecular dynamics modelled by the Boltzmann equation. On one hand, from a detailed microscopic description of the dynamics, relevant scales for problem variables are identified, including, e.g., time, space, and densities. On the other hand, macroscopic, observational variables are compared with typical microscopic variables, which are assumed to be comparatively small. This allows the derivation of simpler, but still precise, models with a focus on observable variables. See [Cer69, Cer88, SR09, GL05, Gol05] and references therein for further information.
In this work, contrary to the usual in the literature, we will call the partial differential equations model a mesoscopic model, and the macroscopic limit will be reserved for the ordinary differential equation model, namely the SIR model, as it is our final goal.
There are several differences between the model introduced here and the one presented in [BDP21] that are worth a brief comment. The most important one is that we include class transitions directly inspired by the SIRS model, with the aim of having as a macroscopic limit, when the problem is spatially homogeneous, the system (1). A second important difference is to consider more general movements of the individuals and proceed to a careful scale separation of the dependent and independent relevant variables and the class transitions. Finally, we study the existence and uniqueness of solutions of the kinetic model, although in a limited manner.
We finish this introduction with an outline of the present work. In Sec. 2 we introduce our basic model, i.e, the kinetic model for epidemiology; in Sec. 3 we formally show that its form macroscopic limit is, in the spatially homogeneous case, the SIRS model (1), validating the idea that the kinetic model is a generalisation of the SIRS model. Furthermore, we define a mesoscopic model, a partial differential equation (PDE) of drift-diffusion-advection type, that lies in between the kinetic and the SIRS model. In Sec.  4, we proceed to show the existence and uniqueness of solutions in the appropriate spaces of particular instances of the kinetic and, by extension, of the PDE model derived before. We show some applications in Sec. 5 and discuss possible limitations and continuations of the present work in Sec. 6.
@@ 2 The Model
Let $f_{j}:\mathbb{R}_{+}\times\Omega\times V$, $j=S,I,R$, be the density of individuals of class $j$ at time $t\geq 0$, position $x\in\Omega\subset\mathbb{R}^{n}$ and velocity $v\in V\subset\mathbb{R}^{n}$. We assume that the space domain $\Omega$ is a bounded open set with smooth boundary and $V$ is compact and spherically symmetrical (in particular, $\int_{V}v\mathrm{d}v=0$). We define $|V|:=\int_{V}\mathrm{d}v$. We will consider that $\Omega$ and $V$ are the same for all individuals except in Subsec. 5.4, where the confining of I individuals in a one-dimensional spatial model is considered. In this case, $\Omega$ is possibly unbounded and $V_{S}$, $V_{R}$, $V_{I}$ are different discrete sets of admissible velocities for S, I, and R individuals, respectively.
@@
In the sequel, whenever we refer to results valid to all classes of individuals, we use the subindex $j$, without further specifications. Furthermore, we use the notation $f_{j}:=f_{j}(t,x,v)$, and $f^{\prime}_{j}:=f_{j}(t,x,v^{\prime})$.
The transition kernel $\mathfrak{Q}_{j}$ describes three different behaviors for the individual in the population: i) a regression toward the mean change of velocity in the space $V$; ii) a preferred direction in which individuals tends to move towards more favorable environments; and iii) a class-transition kernel accounting to the pathogen spread dynamics, healing and immunity waning, according to the SIRS model.
@@
  • The regression toward the mean reorientation kernel is given by
@@
    The reorientation rate is given by $\lambda_{j}>0$. This expression describes a mean reversion process, indicating a drift towards its mean function. In particular, given a point $x\in\Omega$, the number of individuals with velocity $v$ tends to increase if the density function $f_{j}$ is below average and decrease otherwise. We note that the kernel $\mathcal{Q}_{j}^{(R)}$ is conservative, i.e., $\int_{V}\mathcal{Q}_{j}^{(R)}[\mathbf{f}]\mathrm{d}v=0$. This simple fact will be useful in the subsequent calculations.
@@
    $\mathcal{Q}_{S}^{(P)}[\mathbf{f}]:=\int_{V}\bigl(T[f_{I}](v,v^{\prime})f_{S}(t,x,v^{\prime})-T[f_{I}](v^{\prime},v)f_{S}(t,x,v)\bigr)\mathrm{d}v^{\prime},$
    where the transition from $v^{\prime}$ to $v$ is given by
@@
    with $\eta,\xi\geq 0$. In particular, this means that susceptible individuals try to avoid directions with increasing concentrations of I individuals (first term in the right-hand side) and to keep the velocity if it points toward a decrease in the concentration of I individuals. For completeness, we also define preferred direction kernels for I and R classes to be identically null:
    $\mathcal{Q}_{I}^{(P)}[\mathbf{f}]:=\mathcal{Q}_{R}^{(P)}[\mathbf{f}]:=0\ ,\quad\forall\kern 5.0pt\mathbf{f}\ .$
@@
    Class transitions kernels $\mathcal{Q}^{(T)}_{j}$ do not satisfy the conservative property.
@@ 3 The macroscopic limit
In this section, we study the macroscopic limit of the kinetic model in three steps. In Subsec. 3.1 we rewrite Eq. (2) in dimensionless form, and identify small parameters; in the sequel, Subsec. 3.2, we identify an advection-reaction-diffusion partial differential equation (PDE) for the evolution of the density of individuals in the domain $\Omega$. We conclude in Subsec. 3.3, that spatially homogeneous solutions of the previous system of PDEs, whenever they exist, correspond to the solution of the SIRS epidemic model (1).
@@ 3.1 Adimensionalization of the kinetic model
In order to provide the dimensionless form of Eq. (2), we define typical observation time, modulus of displacement, modulus of velocity, and density $t_{0}$, $x_{0}$, $v_{0}$, $f_{0j}$, respectively. See Fig. 2.
We define new dimensionless variables $\tilde{t}:=t/t_{0}$, $\tilde{x}:=x/x_{0}$, $\tilde{v}:=v/v_{0}$, and $\tilde{f}_{j}:=f_{j}/f_{0j}$ such that
@@
Figure 2: Example of a movement of a particle from point A (red) to point B (blue). The particle travels at instantaneous velocity $v$ with trajectory given by the grey line, with typical modulus $v_{0}$. The direction of movement changes periodically, according to the given kernels. After an observation time $t_{0}$, the typical displacement (black arrow) is given by $x$, with modulus $x_{0}$, resulting in a average (macroscopic) velocity of $\frac{x_{0}}{t_{0}}\ll v_{0}$. Typical densities are not represented in the figure, which focus on the movement of a single particle.
We assume that typical density is low (an assumption akin to the rarefied gas assumption used in taking macroscopic limits of ideal gases modeled by Boltzmann equations, cf. [Cer69, Cer88, SR09] and references therein). More precisely, we redefine the transition kernels $\varepsilon^{-2}\tilde{\mathfrak{Q}}_{j}^{(\varepsilon)}[\mathbf{f}]:=\frac{t_{0}}{f_{0j}}\mathfrak{Q}_{j}[\mathbf{f}_{0}\tilde{\mathbf{f}}$], and omit the $\tilde{}$ notation to find the non-dimensional form of the kinetic equation
$\partial_{t}f_{j}+\frac{1}{\varepsilon}v\cdot\nabla_{x}f_{j}=\frac{1}{\varepsilon^{2}}\mathfrak{Q}_{j}^{(\varepsilon)}[\mathbf{f}],$ (3)
Solution of Eq. (3) will be denoted by $f_{j}^{(\varepsilon)}$ to stress its dependency on the small positive dimensionless parameter $\varepsilon$.
@@
$\mathfrak{Q}_{j}^{(\varepsilon)}=\mathcal{Q}_{j}^{(R)}+\varepsilon\mathcal{Q}_{j}^{(P)}+\varepsilon^{2}\mathcal{Q}_{j}^{(T)}+o\left(\varepsilon^{2}\right)\ .$ (4)
This means, in particular, that the leading order effect in the individual movement is the reorientation, followed by a small-order perturbation that gives preferred directions (when it is non-zero), and a second-order effect guaranteeing that the class transitions are compatible with the SIRS model.
@@ 3.2 The mesoscopic limit
In this subsection, we derive a partial differential equation for the space and time dependence of the leading order term, i.e., to
$\rho_{j}:=\lim_{\varepsilon\to 0}\int_{V}f_{j}^{(\varepsilon)}\mathrm{d}v\ ,\quad(t,x)\in\mathbb{R}_{+}\times\Omega\ .$
The model satisfied by $\rho_{j}$ will be called mesoscopic model, which is an intermediate model between the phase-space evolution described by the kinetic equations in $\mathbb{R}_{+}\times\Omega\times V$ and the SIRS model in which all variables depend only on $t\in\mathbb{R}_{+}$.
@@
Theorem 1 (Lax-Milgram Theorem [Ram80, Sho97, BCST26]).
Let $\mathcal{H}$ be a Hilbert space with inner product $\langle\cdot,\cdot\rangle$, and let $\mathcal{B}$ be a continuous bilinear form $\mathcal{B}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb${R}. Assume further that $\mathcal{B}$ is coercive, i.e. $\mathcal{B}(x,x)\geq c\langle x,x\rangle$ for some positive constant $c$. Then, given $w\in\mathcal{H}$, there exists a unique element $x\in\mathcal{H}$ such that $\mathcal{B}(x,u)=\langle w,u\rangle$ for all $u\in\mathcal{H}$.
In the sequel, we will use a more specific form of the Lax-Milgram theorem:
@@
Let $\mathcal{H}_{0}=\{h\in L^{2}(V)|\int_{V}h\mathrm{d}v=0\}$. Then, for each $\mathbf{g}=(g_{S},g_{I},g_{R})$, with $g_{j}\in\mathcal{H}_{0}$ there is a unique $\mathbf{f}=(f_{S},f_{I},f_{R})$ such that $f_{j}\in\mathcal{H}_{0}$, and $\mathcal{Q}^{(R)}_{j}[\mathbf{f}]=-\mathbf{g}$, $j=S,I,R$, in the weak sense.
@@
We start by defining
$W_{\lambda}[f]:=\lambda\left(\frac{1}{|V|}\int_{V}f^{\prime}\mathrm{d}v^{\prime}-f\right)\ ,$
for strictly positive $\lambda$. It is clear that $\mathcal{H}_{0}$ is a Hilbert space with the natural $L^{2}$-inner product, $\langle f,g\rangle:=\int_{V}fg\mathrm{d}v$. Define the bilinear form $\mathcal{B}(g,h):=-\int_{V}W_{\lambda}[g]h\mathrm{d}v$, and
$\mathcal{B}(g,g)=-\lambda\left(\int_{V\times V}\frac{1}{|V|}g^{\prime}g\mathrm{d}v^{\prime}\mathrm{d}v-\int_{V}g^{2}\mathrm{d}v\right)=\lambda\langle g,g\rangle.$
We conclude that $\mathcal{B}$ is coercive. The continuity is immediate.
@@
$\langle-g,h\rangle=-\mathcal{B}(f,h)=\int_{V}W_{\lambda}[f]h\mathrm{d}v\ ,\quad\forall h\in\mathcal{H}_{0}\ .$
Noting that $\mathcal{Q}^{(R)}_{j}[\mathbf{f}]=W_{\lambda_{j}}[f_{j}]$, we finish the proof. ∎
@@
$f_{j}^{(\varepsilon)}=f_{j,0}+\varepsilon f_{j,1}+\varepsilon^{2}f_{j,2}+o\left(\varepsilon^{2}\right)\ ,$
where $f_{j,i}$ is the $i$-th order expansion in $\varepsilon$ of $f_{j}^{(\varepsilon)}$.
The objective of this subsection is to find a system of PDEs for $\rho_{j}=\int_{V}f_{j,0}\mathrm{d}v$. From Eq. (4), it is clear that $\mathcal{Q}^{(R)}_{j}$, $\mathcal{Q}^{(P)}_{j}$, and $\mathcal{Q}^{(T)}_{j}$ should be expanded up to second, first, and zero order in $\varepsilon$, respectively. Ommitting terms of higher order, we find:
$\mathcal{Q}^{(R)}_{j}[\mathbf{f}]$ $=\lambda_{j}\left[\frac{1}{V}\int_{V}\left(f_{j,0}^{\prime}+\varepsilon f_{j,1}^{\prime}+\varepsilon^{2}f_{j,2}^{\prime}\right)\mathrm{d}v^{\prime}-\left(f_{j,0}+\varepsilon f_{j,1}+\varepsilon^{2}f_{j,2}\right)\right]$
$=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{0}]+\varepsilon\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{1}]+\varepsilon^{2}\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{2}]\ ,$
$\mathcal{Q}^{(P)}_{S}[\mathbf{f}]$ $=\int_{V}\left\{T\left[f_{I,0}+\varepsilon f_{I,1}\right](v,v^{\prime})\left(f_{S,0}^{\prime}+\varepsilon f_{S,1}^{\prime}\right)\right$.
$\quad-\left.T\left[f_{I,0}+\varepsilon f_{I,1}\right](v^{\prime},v)\left(f_{S,0}+\varepsilon f_{S,1}\right)\right\}\mathrm{d}v^{\prime}$
$=\int_{V}\left(T[f_{I,0}](v,v^{\prime})f_{S,0}^{\prime}-T[f_{I,0}](v^{\prime},v)f_{S,0}\right)\mathrm{d}v^{\prime}+$
$\quad+\varepsilon\left\{\int_{V}\left(T[f_{I,0}](v,v^{\prime})f_{S,1}^{\prime}-T[f_{I,0}](v^{\prime},v)f_{S,1}\right)\mathrm{d}v^{\prime}\right$.
$\qquad\left.+\int_{V}\left(T[f_{I,1}](v,v^{\prime})f_{S,0}^{\prime}-T[f_{I,1}](v^{\prime},v)f_{S,0}\right)\mathrm{d}v^{\prime}\right\}$
$=\mathcal{Q}^{(P)}_{S}[\mathbf{f}_{0}]+\overline{\mathcal{Q}}^{(P)}_{S}[\mathbf{f}_{0},\mathbf{f}_{1}]\ ,$
where
$\overline{\mathcal{Q}}^{(P)}_{S}[\mathbf{f}_{0},\mathbf{f}_{1}]$ $:=\int_{V}\left(T[f_{I,0}](v,v^{\prime})f_{S,1}^{\prime}-T[f_{I,0}](v^{\prime},v)f_{S,1}\right)\mathrm{d}v^{\prime}$
$\quad+\int_{V}\left(T[f_{I,1}](v,v^{\prime})f_{S,0}^{\prime}-T[f_{I,1}](v^{\prime},v)f_{S,0}\right)\mathrm{d}v^{\prime}$
is conservative and $\mathbf{f}_{\kappa}:=(f_{S,\kappa},f_{I,\kappa},f_{R,\kappa})$, $\kappa=0,1$.
Finally, expanding both sides of Eq. (3), we find
@@
$\qquad+\varepsilon^{2}\left(\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{2}]+\overline{\mathcal{Q}}_{j}^{(P)}[\mathbf{f}_{0},\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(T)}[\mathbf{f}_{0}]\right)\ .$
where $\overline{\mathcal{Q}}_{j}^{(P)}\equiv 0$ for $j=I,R$.